п JCUIMIect.itm m Project Saved Help Save & Exit Se Check my w The Walton Toy Company manufactures a line of dolls and a sewi need all 3 requirements
Answer #1

Solution

Walton Toy Company

1. determination of the number of direct labor hours needed to manufacture one unit of each of the five products:

Product

Direct labor

Direct labor cost per hour

direct labor hours per unit

Debbie

$6.40

$16

0.40

Trish

$4.00

$16

0.25

Sarah

$11.20

$16

0.70

Mike

$8.00

$16

0.50

Sewing Kit

$3.20

$16

0.20

Note:

Direct labor hour per unit is calculated as follows,

Total direct labor hour per unit/direct labor hour

Direct labor hour rate = $16 per hour

For instance, direct labor hour for the product Debbie = $6.40/$16 =0.40 hours

2   determination of variable overhead cost for one unit of each of the five products:

Product

direct labor hours per unit

Variable overhead per hour

Variable overhead per unit

Debbie

0.4

$2

$0.80

Trish

0.25

$2

$0.50

Sarah

0.7

$2

$1.40

Mike

0.5

$2

$1.00

Sewing Kit

0.2

$2

$0.40

Note: variable overhead per unit = direct labor hours per unit x variable overhead rate per hour

For instance, Debbie variable overhead per unit = 0.4 hours x $2 = $0.80

  1. Contribution margin per hour for each of the company’s five products:
  1. Determination of contribution margin per unit and per direct labor hour for all the products:

Product

selling price per unit

Direct material cost per unit

Direct labor cost per unit

Variable overhead per unit

Total variable cost per unit

Contribution margin per unit

direct labor hours per unit

contribution per direct labor hour

Debbie

$16.70

$4.30

$6.40

$0.80

$11.50

$5.20

0.40 hours

$13

Trish

$7.50

$1.10

$4.00

$0.50

$5.60

$1.90

0.25 hours

$7.60

Sarah

$26.60

$6.44

$11.20

$1.40

$19.04

$7.56

0.70 hours

$10.80

Mike

$14.00

$2.00

$8.00

$1.00

$11.00

$3.00

0.50 hours

$6.00

Sewing Kit

$9.60

$3.20

$3.20

$0.40

$6.80

$2.80

0.20 hours

$14.00

Contribution margin per unit = selling price – total variable cost per unit

Total variable cost per unit = direct materials per unit + direct labor per unit + variable overhead per unit

For instance, Contribution margin per unit for Trish is computed as follows,

Total variable cost per unit = 1.10 + 4 + 0.50 = $5.60

Contribution margin per unit = $7.50 - $5.60 = $1.90

Now, contribution margin per hour = contribution margin per unit/hours per unit

For trish,

Contribution margin per hour = $1.90/0.25 hours = $7.60 per hour